| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 2025 121 05 30 50 Motor #2: appears stuck with a reference sensor activated. | 2 | 1 | B12225.224 | TORON | #224 | 05:22:14 | 1 | -1 | |
| rl: | 2025 122 12 16 46 New value set for watermark SUPPLY.VOLTAGE.HIGH | 5 | 3 | B12225.224 | TORON | #224 | 05:22:19 | 1 | -1 |
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 2025 121 05 29 28 Reset Warm -> Tepid from 2025 121 05 29 22 until 2025 121 05 29 23 . Counts: 1,0 | 1 | 3 | B12225.224 | TORON | #224 | 05:22:11 | 1 | -1 | |
| ds: | Suppressed oscillation of filter wheel. | 2 | 5 | B12225.224 | TORON | #224 | 15:19:11 | 1 | -1 |
| _jday | _loc | _serial | _skc | ds | dt | ed | fz | hg | rs | sl | uv | zs |
|---|---|---|---|---|---|---|---|---|---|---|---|---|
| 122 | TORONTO | #224 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 94.275) | 76 | 2 | 2 | 6 | 18 | 2 | 6 | 32 | 41 |
| serial | days |
|---|---|
| 224 | more than a year |
| Serial | last data | days ago |
|---|---|---|
| 224 | 2025123 | 0-1 |
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
|---|---|---|---|---|---|---|---|---|---|
| 224 | 16 | 0 | 2 | 0 | 0 | 0 | 0 | 0 | 0 |
| Serial | min | max | filter | fixable | 30-day history |
|---|---|---|---|---|---|
| #224 | 2668360 | 2669890 | 0 | yes |